# Tora: Trajectory-oriented Diffusion Transformer for Video Generation

### v1

@@ Abstract
Recent advancements in Diffusion Transformer (DiT) have demonstrated remarkable proficiency in producing high-quality video content. Nonetheless, the potential of transformer-based diffusion models for effectively generating videos with controllable motion remains an area of limited exploration. This paper introduces Tora, the first trajectory-oriented DiT framework that integrates textual, visual, and trajectory conditions concurrently for video generation. Specifically, Tora consists of a Trajectory Extractor (TE), a Spatial-Temporal DiT, and a Motion-guidance Fuser (MGF). The TE encodes arbitrary trajectories into hierarchical spacetime motion patches with a 3D video compression network. The MGF integrates the motion patches into the DiT blocks to generate consistent videos following trajectories. Our design aligns seamlessly with DiT’s scalability, allowing precise control of video content’s dynamics with diverse durations, aspect ratios, and resolutions. Extensive experiments demonstrate Tora’s excellence in achieving high motion fidelity, while also meticulously simulating the movement of the physical world.
Figure 1: More generated samples. Tora accommodates a wide array of visual conditions, encompassing both single starting frames and combinations of initial and final frames (as seen in the sixth row), and expertly handles multiple trajectories to precisely manipulate multiple objects. It adeptly facilitates video generation across diverse aspect ratios, resolutions, and durations, thereby ensuring flexible and adaptive content creation. Check out our project page for all video demos.
@@ Introduction
Video generation entails the creation of consistent motion within a sequence of generated images, thereby emphasizing the importance of motion control. Prior research endeavors, such as VideoComposer (Wang et al. 2023a) and DragNUWA (Yin et al. 2023), have successfully implemented generalized motion manipulation techniques. These methods typically involve extracting motion vectors and optical flow data to direct the video generation models. Building upon these advancements, MotionCtrl (Wang et al. 2023b) innovates by separately managing camera movements and object motions, thereby expanding the achievable motion ranges and patterns. Despite promising controllable motion quality, these methods are limited to generating videos of up to 16 frames in length and at a lower, fixed resolution. Moreover, limitations in the number of frames can hinder the seamless portrayal of motion, particularly in scenarios involving significant positional shifts. This often leads to issues such as motion blur, appearance distortions, and unnatural movements, such as parallel drifting, which deviate from real-world physical dynamics.
@@
To address these challenges, we introduce Tora—a pioneering DiT model that simultaneously integrates text, image, and trajectory data, ensuring meticulous and versatile control over video content. Tora innovatively embeds the provided trajectory into multiple patches, enhancing motion fluidity and more closely adhering to the natural principles of object kinetics, leveraging its remarkable scaling properties. Notably, our work adopts OpenSora (Zheng et al. 2024), an open-source version of Sora, as the foundational DiT framework. To facilitate motion control, we design two innovative modules: the Trajectory Extractor (TE), which transforms arbitrary trajectories into hierarchical spacetime motion patches, and the Motion-guidance Fuser (MGF), meticulously crafted to integrate these motion patches within the stacked architecture of DiT blocks. More specifically, TE initially converts positional displacements along trajectories between successive frames into the RGB domain via flow visualization techniques. These visualized displacements undergo Gaussian filtering to mitigate scattered issues. Subsequently, a 3D Variational Autoencoder (VAE) (Kingma and Welling 2013) encodes trajectory visualizations into spacetime motion latents, which share the same latent space with video patches. The motion latents are then decomposed into multiple levels of motion conditions via a series of stacked lightweight modules. Our VAE architecture is inspired by MAGVIT-v2 (Yu et al. 2023b) but simplified by foregoing codebook dependencies and utilizing solely VAE reconstruction loss for training. The MGF integrates adaptive normalization layers (Perez et al. 2018) to infuse multi-level motion conditions into the corresponding DiT blocks. We explored various adaptations of transformer blocks including adaptive layer normalization, cross-attention, and extra channel connections to inject the motion conditions. Among these, adaptive layer normalization emerged as the most effective to generate consistent videos following trajectories.
@@
To train Tora, annotated videos with captions and movement trajectories are essential. We adapt OpenSora’s workflow to transform raw videos into high-quality video-text pairs and leverage an optical flow estimator (Xu et al. 2023) for trajectory extraction. Significantly, we combine motion segmentation results (Zhao et al. 2022) with flow scores to filter out instances that primarily contain camera movement. This strategic approach improves our ability to accurately follow the trajectory of specific objects within videos. As a result, this careful selection process leads to the creation of a dataset containing 630k high-quality video clips with consistent motion. With an adapter-like strategy (Mou et al. 2024), we solely train the temporal blocks, together with the TE and MGF. This strategy seamlessly integrates DiT’s inherent generative knowledge with external motion signals.
@@
  • We introduce Tora, the first trajectory-oriented DiT for video generation. As illustrated in Figure 1, Tora seamlessly integrates a broad range of visual and trajectory instructions, thereby proficiently enabling the creation of motion-manipulable videos.
  • To align with the scalability of DiT, we design a novel trajectory extractor and a motion-guidance fusion mechanism to obtain spacetime motion patches, subsequently injecting these patches into DiT blocks. We ablate several architecture choices and offer empirical baselines for future research in DiT-based motion control.
@@
Diffusion models have demonstrated an impressive capability to generate high-quality video samples. Previous research (Ho et al. 2022b, a; Singer et al. 2022; Wu et al. 2023; Khachatryan et al. 2023; Zhang et al. 2023b) commonly used video diffusion models (VDMs) that incorporated temporal convolutional and attention layers into the pre-trained image diffusion models. Subsequently, VideoCrafter (Chen et al. 2023) and SVD (Blattmann et al. 2023) expand the application of video diffusion models to larger datasets, while TF-T2V (Wang et al. 2024a) directly learn from extensive text-free videos. Lumiere (Bar-Tal et al. 2024) introduces the concept of temporal downsampling within the 3D U-Net framework. Nonetheless, these methods encounter limitations in generating long videos, owing to the inherent constraints on capacity and scalability within the UNet design. Conversely, DiT-based models (Brooks et al. 2024; Zheng et al. 2024; Bao et al. 2024) can directly generate videos extending up to tens of seconds. Sora (Brooks et al. 2024) converts visual data into a unified representation, facilitating large-scale training and enabling the generation of 1-minute high-definition video. Vidu (Bao et al. 2024) is capable of generating both realistic and imaginative videos in various aspect ratios and resolutions. For our study, we adopt OpenSora (Zheng et al. 2024) as the foundational model, which is an open-source alternative to Sora.
Figure 2: Overview of the Tora Architecture. For achieving trajectory-controlled DiT-based video generation, we introduce two novel modules: the Trajectory Extractor and the Motion-guidance Fuser. The Trajectory Extractor employs a 3D motion VAE to embed trajectory vectors into the same latent space as video patches, effectively preserving motion information across consecutive frames. Subsequently, it uses stacked convolutional layers to extract hierarchical motion features. The Motion-guidance Fuser utilizes adaptive normalization layers to seamlessly inject these multi-level motion conditions into the corresponding DiT blocks, ensuring the generation of videos that consistently follow the defined trajectories. Our method aligns with the scalability of DiT, enabling the creation of high-resolution, motion-controllable videos with prolonged durations.
@@
To better control motion in generated video, a multitude of studies have endeavored to introduce diverse motion signals in VDMs. Pioneering works like MotionDirector (Zhao et al. 2023) and VMC (Jeong, Park, and Ye 2024) have utilized reference videos to extract motion patterns applicable to diverse video generations. VideoComposer (Wang et al. 2023a) expands upon this by adopting depth maps, sketches, or motion vectors from references as conditional inputs for motion control. Nonetheless, these methodologies are limited to reproducing existing motion patterns. Conversely, approaches that leverage trajectories or bounding boxes (Yin et al. 2023; Dai et al. 2023; Wang et al. 2024b) in video generation promise greater adaptability and user accessibility. DragNUWA (Yin et al. 2023) breaks new ground by integrating trajectory-based conditioning into VDMs, facilitating complex camera and object movements. AnimateAnything (Dai et al. 2023) employs motion masks for precise control over the moving regions. TrailBlazer (Wan-Duo Kurt Ma 2023), employs explicit attention mechanisms to maneuver generated objects along precise trajectories. MotionCtrl (Wang et al. 2024b) facilitates more flexible control, allowing separate adjustment of both camera movements and individual object motions. However, all of them often yield noticeable artifacts in both motion consistency and visual presentation when applied to longer video sequences. In contrast, our method first integrates trajectories into DiT architecture, specifically designed to accommodate scaling properties. This enables motion fluidity and closer adherence to the physical world.
@@
Diffusion Transformer (DiT). The DiT (Peebles and Xie 2023) constitutes a seminal architectural innovation that synergistically combines the advantages of diffusion models with the transformer architectures (Vaswani et al. 2017). This integration is purposed to transcend the limitations of traditional U-Net-based LDMs, augmenting their performance, versatility, and scalability. While maintaining the overarching formulation congruent with established LDM frameworks, the paradigm shift resides in the substitution of the U-Net with a transformer architecture for the denoising function $\epsilon_{\theta}(\cdot)$ learning, thereby marking a pivotal advance in the realm of generative modeling.
@@
Tora adopts OpenSora as the foundational model for its DiT architecture. To facilitate precise and user-friendly motion control across diverse durations, Tora introduces two novel motion-processing components: the Trajectory Extractor (TE) and the Motion-guidance Fuser (MGF). These modules serve to encode provided trajectories into multi-level spacetime motion patches, meticulously integrating these patches within the stacked structure of DiT blocks. An overview of Tora’s workflow is depicted in Figure 2.
@@
During processing, a video autoencoder (Yu et al. 2023a) is first employed to diminish both spatial and temporal dimensions of videos. To elaborate, it encodes the input video $X\in\mathbb{R}^{L\times H\times W\times 3}$ into video latent $z_{0}\in\mathbb{R}^{l\times h\times w\times 4}$, where $L$ denotes the video length and $l=L/4,h=H/8,w=W/8$. $z_{0}$ is next “patchified”, resulting in a sequence of input tokens $I\in\mathbb{R}^{l\times s\times d}$. Here, $s=hw/p^{2}$ and $p$ denotes the patch size. $I$ is then forwarded to the ST-DiT, which models these compressed representations. In both SSA and TSA, standard Attention is performed using Query (Q), Key (K), and Value (V) matrices:
$Q=W_{Q}\cdot I_{norm};K=W_{K}\cdot I_{norm};V=W_{V}\cdot I_{norm}$ (3)
@@
Here, $I_{norm}$ represents the normalized $I$, $W_{Q},W_{K},W_{V}$ are learnable projection matrices. The textual prompt is first embedded utilizing a T5 text encoder. The cross-attention mechanism is employed between the embedded prompt and the intermediate features from either the SSA or the TSA.
@@
Trajectory Extractor. Trajectories have proven to be a more user-friendly method for controlling the motion of generated videos. Specifically, given a trajectory $traj=\left\{(x_{i},y_{i})\right\}_{i=0}^{L-1}$, where $(x_{i},y_{i})$ denotes the spatial position $(x,y)$ at the $i$-$th$ frame the trajectory passes through. Previous studies primarily encode the horizontal offset $u(x_{i},y_{i})$ and the vertical offset $v(x_{i},y_{i})$ as the motion condition for $(x_{i},y_{i})$:
$u(x_{i},y_{i})=x_{i+1}-x_{i};~{}v(x_{i},y_{i})=y_{i+1}-y_{i}.$ (4)
@@
However, the DiT model employs a video autoencoder and a patchification process to convert the video into video patches. Here, each patch is derived across multiple frames, rendering it inappropriate to straightforwardly employ frame-to-frame offsets. To address this, our TE converts the trajectory into motion patches, which inhabit the same latent space as the video patches. Particularly, we begin by transforming the $traj$ into a trajectory map $g\in\mathbb{R}^{L\times H\times W\times 2}$, enhanced with a Gaussian Filter to mitigate scatter. Notably, the first frame employs a fully-zero map. Afterward, the trajectory map $g$ is translated into the RGB color space, producing $g_{vis}\in\mathbb{R}^{L\times H\times W\times 3}$ through a flow visualization technique. We use a 3D VAE to compress trajectory maps, achieving an 8x spatial and 4x temporal reduction, aligning with OpenSora framework. Our 3D VAE is based on the Magvit-v2 architecture, with spatial compression initialized using the VAE of SDXL (Podell et al. 2023) to accelerate convergence. We train the model using only reconstruction loss to obtain the compact motion latent representation $g_{m}\in\mathbb{R}^{l\times h\times w\times 4}$ from the $g_{vis}$.
@@
To match the size of the video patches, we use the same patch size on $g_{m}$ and encode it using a series of convolutional layers, resulting in spacetime motion patches $f\in\mathbb{R}^{l\times s\times d_{v}}$. Here $d_{v}$ is the dimension of motion patches. The output of each convolutional layer is skip-connected to the input of the next layer to extract multi-level motion features:
$f_{i}=Conv^{i}~{}(f_{i-1})+f_{i-1}.$ (5)
@@
    $h_{i}=\gamma_{i}\cdot h_{i-1}+\beta_{i}+h_{i-1}.$ (7)
@@
    $h_{i}=CrossAttn~{}([h_{i-1},f_{i}])+h_{i-1}.$ (8)
We experiment with three types of fusion architectures, finding that the adaptive norm demonstrates the best generation performance and compute efficiency. For the remainder of the paper, MGF employs the adaptive norm layer unless otherwise specified.
@@
Training Strategy and Data Processing
To achieve fine-grained control while generating a video using arbitrary trajectories, as well as text, images, or their composites, we introduce several training strategies for different condition injections.
Motion condition training. Inspired by DragNUWA and MotionCtrl, we use a two-stage training approach for trajectory learning. In the first stage, we extract the dense optical flow (Xu et al. 2023) from the training video as the trajectory, offering richer information to accelerate motion learning. During the second stage, to adapt the model from complete optical flow to more user-friendly trajectories, we select a random sample of 1 to $N$ object trajectories from the optical flow according to the motion segmentaton results and flow scores. To address the scattered issues of sparse trajectories, we apply a Gaussian filter for refinement. Upon completion of the two-stage training phase, Tora enables flexible motion control using arbitrary trajectories.
@@
Data Processing. For Tora, it is crucial that the video clips in the training dataset are annotated with both captions and movement trajectories (optical flow). To fulfill this requirement, we employ a structured data processing method. Initially, raw videos are segmented into shorter clips based on scene detection. Subsequently, these clips are evaluated using aesthetic scores. Given our criteria, static videos are excluded by retaining only those with optical flow scores exceeding 3. To better obtain object trajectories, We employ the results of motion segmentation and camera detection to remove instances that primarily contain camera trajectories in dynamic scenes. Additionally, dramatic camera or object motions in some videos can cause significant optical flow deviations, interfering with trajectory training. Consequently, we retain these videos with a probability of $(1-score/100)$. For eligible videos, we generate captions using the PLLaVA model (Xu et al. 2024). To balance speed and performance, we utilize the 13B version.
@@
Implementation Details. We train Tora based on OpenSora v1.2 weights. The resolutions of the training videos vary from 144p to 720p. The numbers of frames range from 51 to 204. To balance the training FLOP and required memory for different resolutions and numbers of frames per iteration, we adjust the batch size accordingly from 1 to 25. We use Adam Optimizer (Kingma and Ba 2015) with a learning rate of $2\times 10^{-5}$ on 4 NVIDIA A100. We train Tora with the dense optical flow for 2 epochs and fine-tune it with the sparse optical flow for 1 epoch. The maximum number of sampling trajectories $N$ is set to 16. The inference step and the guidance scale are set to 30 and 7.0, respectively.
Dataset. Our training videos are sourced from three datasets: 1) Panda-70M (Chen et al. 2024), from which we use the training-10M subset containing high-quality videos; 2) Mixkit (Envato 2024); and 3) Interval Video. After applying our data processing pipeline, we select approximately 630k eligible videos for the training dataset. For inference, we curate 185 clips with diverse motion trajectories and scenes, to serve as a new benchmark for evaluating the motion controllability. Additional information about the construction of the evaluation datasets is available in our supplementary materials.
Metrics. We leverage standard metrics such as Fréchet Video Distance (FVD) (Unterthiner et al. 2018), and CLIP Similarity (CLIPSIM) (Wu et al. 2021) to quantitatively evaluate video quality. For assessing motion controllability, we leverage Trajectory Error (TrajError), which computes the average distance between the generated and pre-defined trajectories. Human evaluation is also introduced.

### v3

@@ Abstract
Recent advancements in Diffusion Transformer (DiT) have demonstrated remarkable proficiency in producing high-quality video content. Nonetheless, the potential of transformer-based diffusion models for effectively generating videos with controllable motion remains an area of limited exploration. This paper introduces Tora, the first trajectory-oriented DiT framework that concurrently integrates textual, visual, and trajectory conditions, thereby enabling scalable video generation with effective motion guidance. Specifically, Tora consists of a Trajectory Extractor (TE), a Spatial-Temporal DiT, and a Motion-guidance Fuser (MGF). The TE encodes arbitrary trajectories into hierarchical spacetime motion patches with a 3D video compression network. The MGF integrates the motion patches into the DiT blocks to generate consistent videos that accurately follow designated trajectories. Our design aligns seamlessly with DiT’s scalability, allowing precise control of video content’s dynamics with diverse durations, aspect ratios, and resolutions. Extensive experiments demonstrate Tora’s excellence in achieving high motion fidelity, while also meticulously simulating the intricate movement of the physical world.
@@ Introduction
Figure 1: More generated samples. Tora accommodates diverse conditions such as text-only inputs, single starting frames, and combinations of initial and final frames (as illustrated in the fifth row). It effectively manages multiple trajectories, allowing for precise manipulation of several objects. Furthermore, Tora supports video generation across different aspect ratios, resolutions, and durations, ensuring flexible content creation. For video demonstrations, please refer to our project page.
@@
Video generation requires consistent motion across image sequences, underscoring the importance of motion control. Previous works, such as VideoComposer (Wang et al. 2023) and DragNUWA (Yin et al. 2023), have implemented generalized motion manipulation through motion vectors and trajectories. Building on this foundation, MotionCtrl (Wang et al. 2024b) innovates by independently managing camera and object motions, thereby expanding the diversity of achievable motion patterns. Despite their promising controllable motion quality, U-Net methods are restricted to generating videos of only 16 frames at a fixed, lower resolution. This limitation hinders the smooth portrayal of motion, particularly during significant positional shifts in the provided trajectory, leading to distortion and unnatural movements, such as parallel drifting, which diverge from real-world dynamics. Consequently, there is an urgent need for a model capable of producing longer videos with robust motion control and detailed physical representations.
@@
To address these challenges, we present Tora, the first DiT model that simultaneously integrates text, images, and trajectories, enabling scalable video generation with robust motion control. Notably, our work adopts OpenSora (Zheng et al. 2024), an open-source version of Sora, as the foundational DiT model. To align motion control with the scalability of the DiT framework, we propose two novel modules: the Trajectory Extractor (TE), which converts arbitrary trajectories into hierarchical spacetime motion patches, and the Motion-guidance Fuser (MGF), designed to seamlessly integrate these patches within the stacked DiT blocks. More specifically, TE initially converts positional displacements along trajectory into the RGB domain via flow visualization techniques. These visualized displacements undergo Gaussian filtering to mitigate scattered issues. Subsequently, a 3D Variational Autoencoder (VAE) (Kingma and Welling 2013) encodes trajectory visualizations into spacetime motion latents, which share the same latent space with video patches. The motion latents are then decomposed into multiple levels of motion conditions via stacked lightweight modules. Our VAE architecture is inspired by MAGVIT-v2 (Yu et al. 2023b) but simplified by foregoing codebook dependencies. The MGF integrates adaptive normalization layers (Perez et al. 2018) to infuse multi-level motion conditions into the corresponding DiT blocks. We explored various adaptations of transformer blocks including adaptive layer normalization, cross-attention, and extra channel connections to inject the motion conditions. Among these, adaptive layer normalization emerged as the most effective to generate consistent videos following trajectory.
@@
During training, we adapt OpenSora’s workflow to generate high-quality video-text pairs and utilize an optical flow estimator (Xu et al. 2023) for trajectory extraction. We also integrate a motion segmentor (Zhao et al. 2022) with a camera detector [^1] https://github.com/antiboredom/camera-motion-detector to filter out instances dominated by camera motion, thereby improving our tracking of specific object trajectories. This careful selection process results in a dataset of 630k high-quality videos with consistent motion. With an adapter-like strategy (Mou et al. 2024), we solely train the temporal blocks, together with the TE and MGF. This strategy seamlessly integrates DiT’s inherent generative knowledge with external motion signals.
@@
  • We introduce Tora, the first trajectory-oriented DiT model for scalable video generation with strong motion guidance. As illustrated in Figure 1, Tora seamlessly integrates various text, visual and trajectory instructions, enabling the creation of motion-manipulable videos.
  • We propose a novel trajectory extractor and a motion-guidance fusion mechanism to facilitate motion control that aligns with the scalability of DiT. Additionally, we ablate several architecture choices and offer empirical baselines for future research.
@@
Diffusion models have demonstrated an impressive capability to generate high-quality video samples. Previous research (Ho et al. 2022b, a; Singer et al. 2022; Khachatryan et al. 2023; Zhang et al. 2023b) commonly used video diffusion models (VDMs) that incorporated temporal convolutional and attention layers into the pre-trained image diffusion models. Subsequently, VideoCrafter (Chen et al. 2023) and SVD (Blattmann et al. 2023) expand the application of video diffusion models to larger datasets, while TF-T2V (Wang et al. 2024a) directly learn from extensive text-free videos. Nonetheless, these methods encounter limitations in generating long videos, owing to the inherent constraints on capacity and scalability within the U-Net design. Conversely, DiT-based models (Brooks et al. 2024; Zheng et al. 2024; Bao et al. 2024) can directly generate videos extending up to tens of seconds. Sora (Brooks et al. 2024) converts visual data into a unified representation, facilitating large-scale training and enabling the generation of 1-minute high-definition video. Vidu (Bao et al. 2024) is capable of generating both realistic and imaginative videos in various aspect ratios and resolutions. For our study, we adopt OpenSora (Zheng et al. 2024) as the foundational model, which is an open-source alternative to Sora.
Figure 2: Overview of the Tora Architecture. We introduce two novel modules: the Trajectory Extractor and the Motion-guidance Fuser. The Trajectory Extractor uses a 3D motion VAE to embed trajectory vectors into the same latent space as video patches, preserving motion information across frames. It then employs stacked convolutional layers to extract hierarchical motion features. The Motion-guidance Fuser utilizes adaptive normalization layers to integrate these multi-level motion conditions into the corresponding DiT blocks, ensuring that generated videos consistently follow defined trajectories. Our method leverages the scalability of DiT, enabling the creation of motion-controllable videos of extended duration.
@@
To better control motion in generated video, a multitude of studies have endeavored to introduce diverse motion signals in VDMs. Pioneering works like MotionDirector (Zhao et al. 2023) and VMC (Jeong, Park, and Ye 2024) have utilized reference videos to extract motion patterns applicable to diverse video generations. VideoComposer (Wang et al. 2023) expands upon this by adopting depth maps, sketches, or motion vectors from references as conditional inputs for motion control. Nonetheless, these methodologies are limited to reproducing existing motion patterns. Conversely, approaches that leverage trajectories or bounding boxes (Yin et al. 2023; Dai et al. 2023; Wang et al. 2024b) in video generation promise greater adaptability and user accessibility. DragNUWA (Yin et al. 2023) breaks new ground by integrating trajectory-based conditioning into VDMs, facilitating complex camera and object movements. AnimateAnything (Dai et al. 2023) employs motion masks for precise control over the moving regions. TrailBlazer (Wan-Duo Kurt Ma 2023), employs explicit attention mechanisms to maneuver generated objects along precise trajectories. MotionCtrl (Wang et al. 2024b) facilitates more flexible control, allowing separate adjustment of both camera movements and individual object motions. However, all of them yield noticeable artifacts in both motion consistency and visual presentation when applied to longer sequences. In contrast, our method first integrates trajectories into DiT architecture, specifically designed to accommodate scaling properties. This enables closer adherence to the physical world.
@@
Diffusion Transformer (DiT). The DiT (Peebles and Xie 2023) introduces a novel architecture that merges the strengths of diffusion models with transformer architectures (Vaswani et al. 2017). This integration aims to address the limitations of traditional U-Net-based latent diffusion models (LDMs), improving their performance, versatility, and scalability. While keeping the overall framework consistent with existing LDMs, the key shift lies in replacing the U-Net with a transformer architecture for learning the denoising function $\epsilon_{\theta}(\cdot)$, thereby marking a pivotal advance in the realm of generative modeling.
@@
Tora employs the Spatial-Temporal Diffusion Transformer (ST-DiT) from OpenSora as its foundational model. To facilitate user-friendly motion control while aligning with the scalability of DiT, Tora integrates two novel motion-processing components: the Trajectory Extractor (TE) and the Motion-guidance Fuser (MGF). An overview of Tora’s workflow is illustrated in Figure 2.
@@
During processing, a video autoencoder (Yu et al. 2023a) is first employed to diminish both spatial and temporal dimensions of videos. To elaborate, it encodes the input video $X\in\mathbb{R}^{L\times H\times W\times 3}$ into video latent $z_{0}\in\mathbb{R}^{l\times h\times w\times 4}$, where $L$ denotes the video length and $l=L/4,h=H/8,w=W/8$. $z_{0}$ is next “patchified”, resulting in a sequence of input tokens $I\in\mathbb{R}^{l\times s\times d}$. Here, $s=hw/p^{2}$ and $p$ denotes the patch size. In both SSA and TSA, standard Attention is performed using Query (Q), Key (K), and Value (V) matrices:
$Q=W_{Q}\cdot I_{norm};K=W_{K}\cdot I_{norm};V=W_{V}\cdot I_{norm},$ (3)
@@
Here, $I_{norm}$ is the normalized $I$, $W_{Q},W_{K},W_{V}$ are learnable matrices. The textual prompt is embedded with a T5 encoder and integrated using a cross-attention mechanism.
@@
Trajectory Extractor. Trajectories have proven to be a more user-friendly method for controlling the motion of generated videos. Specifically, given a trajectory $traj=\left\{(x_{i},y_{i})\right\}_{i=0}^{L-1}$, where $(x_{i},y_{i})$ denotes the spatial position $(x,y)$ at the $i$-$th$ frame the trajectory passes through. Previous studies primarily encode the horizontal offset $u(x_{i},y_{i})$ and the vertical offset $v(x_{i},y_{i})$ as the motion condition:
$u(x_{i},y_{i})=x_{i+1}-x_{i};~{}v(x_{i},y_{i})=y_{i+1}-y_{i},$ (4)
@@
However, the DiT model employs a video autoencoder and a patchification process to convert the video into patches. Here, each patch is derived across multiple frames, rendering it inappropriate to straightforwardly employ frame-to-frame offsets. To address this, our TE converts the trajectory into motion patches, which inhabit the same latent space as the video patches. Particularly, we begin by transforming the $traj$ into a trajectory map $g\in\mathbb{R}^{L\times H\times W\times 2}$, enhanced with a Gaussian Filter to mitigate scatter. Notably, the first frame employs a fully-zero map. Afterward, the trajectory map $g$ is translated into the RGB color space, producing $g_{vis}\in\mathbb{R}^{L\times H\times W\times 3}$ through a flow visualization technique. We use a 3D VAE to compress trajectory maps, achieving an 8x spatial and 4x temporal reduction, aligning with OpenSora framework. Our 3D VAE is based on the Magvit-v2 architecture, with spatial compression initialized using the VAE of SDXL (Podell et al. 2023) to accelerate convergence. We train the model using only reconstruction loss to obtain the compact motion latent representation $g_{m}\in\mathbb{R}^{l\times h\times w\times 4}$ from the $g_{vis}$.
@@
To match the size of the video patches, we use the same patch size on $g_{m}$ and encode it using a series of convolutional layers, resulting in spacetime motion patches $f\in\mathbb{R}^{l\times s\times d^{\prime}}$. Here $d^{\prime}$ is the dimension of motion patches. The output of each convolutional layer is skip-connected to the input of the next layer to extract multi-level motion features:
$f_{i}=Conv^{i}~{}(f_{i-1})+f_{i-1},$ (5)
@@
    $h_{i}=\gamma_{i}\cdot h_{i-1}+\beta_{i}+h_{i-1},$ (7)
@@
    $h_{i}=CrossAttn~{}([h_{i-1},f_{i}])+h_{i-1},$ (8)
We evaluate three types of fusion architectures and find that the adaptive norm yields the best performance and computational efficiency. For the remainder of the paper, MGF employs the adaptive norm layer unless otherwise specified.
@@
Data Processing and Training Strategy
Data Processing. We employ a structured data processing method to obtain high-quality training videos with consistent object motion. Initially, raw videos are segmented into shorter clips based on scene detection [^2] https://github.com/Breakthrough/PySceneDetect. Subsequently, we remove invalid videos, such as those with encoding errors, a duration of zero, and low resolution. Furthermore, we utilize aesthetic [^3] https://github.com/christophschuhmann/improved-aesthetic-predictor and optical flow scores (Xu et al. 2023) to filter out low-quality videos. To concentrate on the motion of primary objects, we implement camera motion filtering, using results from motion segmentation (Zhao et al. 2022) and camera detection to exclude instances predominantly exhibiting camera movement. Dramatic object motions in certain videos can lead to significant optical flow deviations, which may interfere with trajectory training. To address this, we retain these videos based on a probability of $(1-score/100)$. For eligible videos, we generate captions using the PLLaVA model (Xu et al. 2024). During inference, we utilize GPT-4o (OpenAI 2023) to refine prompts, ensuring alignment with training process for consistency. More detailed information about the filtering and the refinement process can be found in the supplementary materials.
Motion condition training. Inspired by DragNUWA and MotionCtrl, we adopt a two-stage training approach for trajectory learning. In the first stage, we extract dense optical flow (Xu et al. 2023) from the training video as the trajectory, providing richer information to enhance motion learning. In the second stage, we adjust the model from complete optical flow to more user-friendly trajectories by randomly selecting 1 to $N$ object trajectories based on motion segmentation results and flow scores. To improve the scattered nature of sparse trajectories, we apply a Gaussian filter for refinement. After completing the two-stage training, Tora facilitates flexible motion control using arbitrary trajectories.
@@
Implementation Details. Tora is initialized with OpenSora v1.2 weights, and training videos have resolutions from 144p to 720p and frame counts ranging from 51 to 204. To balance training FLOP and memory usage, we adjust the batch size from 1 to 50. We use Adam Optimizer (Kingma and Ba 2015) with a learning rate of $2\times 10^{-5}$ on 4 NVIDIA A100. The 3D VAE in TE is initially trained on datasets  (Mehl et al. 2023; Mayer et al. 2016; Ranjan et al. 2020; Cabon, Murray, and Humenberger 2020) annotated with optical flow and then frozen during Tora training. We train Tora for 2 epochs with dense optical flow and fine-tune for 1 epoch with sparse trajectories. The maximum number of sampling trajectories $N$ is set to 16. The inference step and the guidance scale are set to 30 and 7.0, respectively.
Dataset. Our training videos are sourced from three datasets: 1) Panda-70M (Chen et al. 2024), from which we use the training-10M subset containing high-quality videos; 2) Mixkit (Envato 2024); and 3) Internal videos. The internal videos are manually annotated, with each clip labeled to include object masks and camera movement. Following our data processing pipeline, we select about 630k eligible videos for the training dataset. For inference, we curate 185 clips with diverse motion trajectories and scenes, to serve as a new benchmark for evaluating the motion controllability.
Metrics. We leverage standard metrics such as Fréchet Video Distance (FVD) (Unterthiner et al. 2018), and CLIP Similarity (CLIPSIM) (Wu et al. 2021) to quantitatively evaluate video quality. For assessing motion controllability, we leverage Trajectory Error (TrajError), which computes the average distance between the generated and pre-defined trajectories. Human evaluation is also introduced in supplementary materials due to space limitations.
Results
We compare our method with popular motion-guided video generation approaches in three settings: 16, 64, and 128 frames, all at a resolution of 512 × 512 for a fair evaluation. The provided trajectories are adjusted to fit the different video lengths. For most U-Net-based methods, we use sequenced inference, where the last frame generated from one batch serves as the visual condition for the next, aligning with their inference strategies. As shown in Table 1, in the 16-frame setting typical for U-Net methods, MotionCtrl and DragNUWA align better with the provided trajectories but still fall short compared to our proposed Tora. With an increasing number of frames, U-Net methods exhibit significant deviations, leading to misalignment errors that result in deformations, motion blur, or object disappearance in later sequences. In contrast, Tora remains highly robust across varying frame counts due to the transformer’s scaling abilities. In the 128-frame test setting, Tora’s trajectory accuracy surpasses other methods by 3 to 5 times, showcasing its outstanding motion control capabilities. Figure 4 presents an analysis of Trajectory Error across different resolutions and durations. Unlike U-Net models, which exhibit substantial trajectory errors over time, Tora shows only a gradual increase in error as duration extends. This aligns with the decrease in video quality observed in the DiT model. The results demonstrate that our method effectively maintains trajectory control over longer durations.